### v1

# Searching for the tetraneutron resonance on the lattice

Abstract
The nature of the tetraneutron ($4n$) system remains a pivotal question in nuclear physics. We investigate the $4n$ system using nuclear lattice effective field theory in finite volumes with a lattice size up to $L=30$ fm, employing both a high-precision N3LO interaction and a simplified SU(4) symmetric one. The ground-state energy is found to decrease smoothly with increasing box size, showing no plateau characteristic of a resonance. We further compute the dineutron-dineutron scattering phase shift using Lüscher’s finite-volume method. The phase shift is negative at low momenta, indicating repulsion in the dilute limit. At intermediate momenta, it exhibits a weak attraction with a peak of approximately $10^{\circ}$ at relative momentum of 60-84 MeV. While this structure does not constitute a resonance, the corresponding confined $4n$ energy of 1.7–3.3 MeV lies close to the experimentally observed low-energy peak.
today
I Introduction
Bound nuclear systems conventionally form when protons and neutrons coalesce through the strong interaction. However, the combination of attractive strong interactions and the absence of Coulomb repulsion makes the existence of pure neutron nuclei a plausible concept, a quest that dates back to the early 1960s Ogloblin and Penionzhkevich (1989). Confirming or ruling out the existence of such neutral nuclei, whether as bound or resonant states, is crucial for advancing our understanding of nuclear physics, the strong interaction, universal fermion properties in the unitary limit Elhatisari et al. (2017); Braaten and Hammer (2006), and astrophysical objects Matsuki et al. (2025); Ivanytskyi et al. (2019).
[…]
While a bound $4n$ nucleus is generally believed to be excluded Lazauskas et al. (2023); Marqués and Carbonell (2021); Bertulani and Zelevinsky (2003), the interpretation of these signals as a resonance remains highly debated. A variety of theoretical approaches, including Jost function analysis in the complex momentum plane Sofianos et al. (1997), the hyperspherical harmonic method Grigorenko et al. (2004), Faddeev-Yakubovsky equations Lazauskas and Carbonell (2005); Hiyama et al. (2016); Carbonell et al. (2017), the Gaussian expansion method Hiyama et al. (2016); Carbonell et al. (2017), the no-core Gamow shell model and density matrix renormalization group Fossez et al. (2017), the $4n$ response function Lazauskas et al. (2017), Alt, Grassberger, and Sandhas (AGS) equations Deltuva (2018); Deltuva and Lazauskas (2019b), and adiabatic hyperspherical methods Higgins et al. (2020, 2021), converge on the absence of a $4n$ resonance. The low-energy peak observed by Duer and others (2022) has been explained as a consequence of dineutron-dineutron correlations through a reaction model study Lazauskas et al. (2023). Reactions with multi-neutron final states using a nonrelativistic conformal field theory are investigated in Hammer and Son (2021). Recent tensor-optimized antisymmetrized molecular dynamics studies suggest that a $4n$ resonance at the nuclear surface may be influenced by core attraction Wan et al. (2025). For further insight into multi-neutron correlations in light nuclei, see Zhang et al. (2025b).
In contrast, many-body methods that employ an external potential trap to confine the $4n$ system, subsequently extrapolating to the no-trap limit, tend to support the existence of a $4n$ resonance with energies and widths close to the experimental values. These include quantum Monte Carlo Pieper (2003); Gandolfi et al. (2017), the no-core shell model Shirokov et al. (2016), and the no-core Gamow shell model Li et al. (2019). However, as discussed in Refs. Deltuva and Lazauskas (2019a); Ishikawa (2020), threshold effects are critical and must be carefully accounted for in such extrapolations. For a comprehensive review of previous efforts, see Marqués and Carbonell (2021).
To further elucidate the nature of the $4n$ signal reported in Duer and others (2022), additional theoretical investigations are needed. In this work, we study the $4n$ system using nuclear lattice effective field theory (NLEFT) Lee (2025); Lähde and Meißner (2019), a method that has proven very successful in providing a unified description of nuclear phenomena on a discretized lattice, including scattering Elhatisari et al. (2025); Wu et al. (2025); Elhatisari et al. (2017, 2015), clustering Epelbaum et al. (2011, 2012); Shen et al. (2023, 2025), nuclear matter Tong et al. (2025a, b); Ma et al. (2024); Ren et al. (2024); Lu et al. (2020), and the structure of light and heavy nuclei Hildenbrand et al. (2025); Ren et al. (2025); Song et al. (2025); Niu and Lu (2025); Wang et al. (2025); Zhang et al. (2025a); Hildenbrand et al. (2024); Meißner et al. (2024); Sarkar et al. (2023); Lu et al. (2022); Shen et al. (2021). We calculate the ground-state energy of $4n$ in finite volumes up to a lattice size of $L\simeq 30$ fm. No external trap is applied, as the finite volume itself provides the necessary confinement for the dilute system.
II Formalism
In nuclear lattice effective field theory, the many-nucleon system is solved on a discretized space with lattice spacing $a$ and $L\times L\times L$ sites. The ground state $|\Psi\rangle$ is obtained by applying Euclidean time projection to an initial wave function $|\Psi_{0}\rangle$ that has nonzero overlap with the true ground state,
$|\Psi\rangle=\lim_{\tau\to\infty}e^{-H\tau}|\Psi_{0}\rangle,$ (1)
[…]
We employ periodic boundary conditions. The finite lattice box confines the nucleons, regardless of whether they form a bound state or a continuum state. To better capture the dilute nature of the $4n$ system and accelerate the convergence to the ground state in Eq. (1), we use a plane-wave Slater determinant as the initial wave function. The single-particle wavefunction is
$\psi_{\mathbf{k}}(\mathbf{n})=\exp\left(i\frac{2\pi}{L}\mathbf{k}\cdot\mathbf{n}\right),$ (2)
with $\mathbf{n}=(n_{x},n_{y},n_{z})$ is the lattice coordinate vector, with each component ranging from 0 to L1L-1, and $\mathbf{k}$ is the lattice momentum vector. The initial state $|\Psi_{0}\rangle$ is constructed as a Slater determinant from the two lowest momentum states, each with spin up and down. Furthermore, we perform an angular momentum projection onto the $A_{1}^{+}$ irreducible representation of the cubic group to project out the $0^{+}$ state and accelerate the Euclidean time evolution Johnson (1982); Lu et al. (2014). The ground state energy at a given projection time $\tau$ is calculated as
$E(\tau)=\frac{\langle\Psi_{0}|e^{-H\tau/2}He^{-H\tau/2}|\Psi_{0}\rangle}{\langle\Psi_{0}|e^{-H\tau}|\Psi_{0}\rangle}.$ (3)
It is well known that the tetraneutron system lacks bound subsystems. However, when confined in a finite volume, the dineutron becomes quasi-bound. Following the method for dimer-dimer scattering on a lattice Elhatisari et al. (2017), we compute the 2n2n-$2n$ scattering phase shifts using Lüscher’s formula. It relates the two-body energy levels in a periodic cubic box to the elastic scattering phase shifts Luscher (1986, 1991):
$p\cot\delta(p)=\frac{1}{\pi L}{S(\eta)},\quad\eta=\left(\frac{pL}{2\pi}\right)^{2},$ (4)
with $S(\eta)$ the regulated three-dimensional zeta function
$S(\eta)=\lim_{\Lambda\to\infty}\left[\sum_{\vec{n},\,|\vec{n}|\leq\Lambda}\frac{1}{\vec{n}^{2}-\eta}-4\pi\Lambda\right].$ (5)
The sum runs over all integer vectors $\mathbf{n}$. The energy of the two-dineutron system in the finite volume, $E^{(L)}$, is related to the relative momentum $p$ by
$E^{(L)}=$ $\frac{p^{2}}{2\mu}-2B_{2n}-2\Delta B_{2n}^{(L)}\left[\sum_{\mathbf{k}}\frac{1}{(\mathbf{k}^{2}-\eta)^{2}}\right]^{-1}$
$\times\sum_{\mathbf{k}}\frac{\sum_{i=1}^{3}\cos(2\pi\alpha k_{i})}{3(\mathbf{k}^{2}-\eta)^{2}}.$ (6)
Here, $\mu$ is the reduced mass of the two dineutrons, $B_{2n}$ is the dineutron binding energy, $\alpha=1/2$ for the dineutron, and $\Delta B_{2n}^{(L)}=B_{2n}^{(L)}-B_{2n}$ the finite volume correction. The last term is a topological factor that accounts for the finite-volume momentum-dependent effects Bour et al. (2011), where the summation over $\mathbf{k}$ includes all integer vectors.
III Results and discussion
We employ a lattice spacing of $a=1.32$ fm, with $L$ ranging from 8 to 23, corresponding to physical box sizes from approximately 10 fm to 30 fm. For comparison, we adopted an interaction with SU(4) symmetry to fit the 1S0 scattering phase shifts. Details regarding its specific form and parameters are provided in the supplementary material.
Figure 1: Ground-state energies of the $4n$ system calculated using NLEFT with both N3LO and SU(4) interactions, and the $2n$ system calculated using the Lanczos eigenvector method with the SU(4) interaction, as functions of the lattice size $L$. The error bars of NLEFT results indicate the uncertainty from the extrapolation in Eq. (S1).
Fig. 1 shows the calculated energy of the tetraneutron system at different $L$ values, using both the full N3LO interaction ($a=1.32$ fm) and the simple SU(4) interaction ($a=1.64$ fm). We also indicate candidate resonance energies from experimental studies: $E=2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}\,$MeV from Ref. Duer and others (2022) and $E=0.83\pm 0.65\text{(stat.)}\pm 1.25\text{(sys.)}\,$MeV from Ref. Kisamori and others (2016). For comparison, the energy of the $2n$ system at different $L$ values using the SU(4) and N3LO interactions is computed via the Lanczos eigenvector method Lanczos (1950), which avoids Monte Carlo and extrapolation uncertainties.
Both the N3LO and SU(4) interactions yield similar results for the $4n$ system, showing strongly repulsive behavior at small $L$ due to Pauli blocking, with energies gradually decreasing as $L$ increases. This consistency confirms earlier findings Lazauskas and Carbonell (2005); Hiyama et al. (2016); Gandolfi et al. (2017); Shirokov et al. (2016); Fossez et al. (2017); Deltuva (2018); Li et al. (2019); Marqués and Carbonell (2021) that the tetraneutron system is insensitive to details of the nuclear interaction as long as it is realistic. The smooth, continuous decrease of the $4n$ energy lacks the characteristic plateau formation expected for resonances in finite-volume methods Hazi and Taylor (1970); Zhang et al. (2008), suggesting a behavior typical of a non-bound system that does not form a resonance. Furthermore, we have examined the energy derivatives and find no evidence of resonant behavior in the tetraneutron system 1.
Recent studies of the $4n$ energy distribution using the 8He($p,p^{4}$He)$4n$ reaction model suggest that the sharp low-energy peak observed by Duer and others (2022) can be explained by dineutron-dineutron correlations Lazauskas et al. (2023). In a finite volume, the $2n$ subsystem becomes quasi-bound in the sense that the lowest $S$-wave level is shifted to negative energy relative to the two-neutron threshold, even though the infinite-volume $nn$ system corresponds to a virtual state. This motivates us to study 2n2n-$2n$ scattering phase shifts using Lüscher’s finite-volume method, following Ref. Elhatisari et al. (2017). In this context, we employ a finite-volume dineutron approximation, which means that the two-neutron subsystem is treated as an effective weakly bound dimer inside the box, characterized by a large scattering length.
Figure 2: Neutron-neutron ($n$-$n$) and neutron-dineutron (nn-$2n$) scattering phase shifts calculated using Lüscher’s finite-volume method with the SU(4) interaction at a lattice spacing a=1.64a=1.64\,fm. Points represent lattice data for different $B_{2n}$ values, and lines show fits based on the effective range expansion.
Before turning to 2n2n-$2n$ scattering, we validate this finite-volume strategy by studying $n$-$n$ and nn-$2n$ scattering. Since at low energies and large separations the interaction details are then suppressed, we consider the zero range leading interaction here. We tune the coupling constant of the zero-range neutron-neutron interaction so that the two-neutron subsystem is close to threshold, and we parameterize this proximity by an effective “binding” scale $B_{2n}$ spanning $0.01$ to $0.09$ MeV. We then use the resulting couplings in the two-neutron and three-neutron systems and compute the finite-volume scattering energy levels, from which we extract the phase shifts using Lüscher’s method, as shown in Fig. 2. Here we define all quantities in lattice units, i.e., physical units multiplied by the corresponding power of $a$.
The lattice data (points) and effective-range-expansion fits (lines) show that $p\cot\delta$ versus $p$ collapses onto the expected universal trend for all systems except at $B_{2n}=0.01$ MeV. This behavior is related to the fact that for short-range interactions with a large two-body scattering length, low-energy observables become insensitive to microscopic details of the interaction and are governed instead by universal large-scattering-length physics associated with the unitarity (conformal) fixed point Braaten and Hammer (2006). In this limit, dimensionless combinations such as $a\,p\cot\delta$ versus apa\,p approach universal functions, and ratios of scattering lengths (e.g., $a_{n\text{-}2n}/a_{n\text{-}n}$) tend toward constants, up to subleading range and regulator effects.
The behavior at $B_{2n}=0.01$ MeV reflects that we are extremely close to the threshold regime where the intrinsic two-neutron length scale becomes very large, and small finite-volume and lattice-spacing effects can be amplified and even shift the two-body pole between a shallow bound state and a virtual state. Consistent with this interpretation, for all lattice spacings examined the neutron–neutron scattering length at $B_{2n}=0.01$ MeV becomes large and negative, indicating a virtual state, and the neutron–dineutron system shows a similarly large and negative scattering length.
Using the effective-range expansion, we extract the ratio of neutron-dineutron to neutron-neutron scattering lengths, $a_{n\text{-}2n}/a_{n\text{-}n}$. The results, summarized in Table 1, show a nearly constant ratio for most cases, in agreement with lattice calculations $1.176(6)$ Elhatisari et al. (2017) and semi-analytic continuum calculations $1.1791(2)$ Bedaque and van Kolck (1998); Bedaque et al. (1998); Gabbiani et al. (2000); Bedaque et al. (2003). We have verified that the results are robust against variations in lattice spacing from $a=1.32$ to $1.97$ fm 1.
Since Lüscher’s relation is derived for elastic scattering, another important consistency check is that the finite-volume levels used to extract nn-$2n$ phase shifts are not significantly affected by the opening of the breakup channel ($n+2n\to n+n+n$). In the finite-volume dimer picture, the relevant inelastic scale is set by the separation energy of the quasi-bound $2n$ level, which we characterize by $B_{2n}$. This corresponds to a characteristic momentum scale $\sqrt{mB_{2n}}$, with $m$ the nucleon mass. For our weakest binding this scale is small, nevertheless, even when the extracted momenta are larger than this estimate, we do not observe systematic departures from the effective-range-expansion description or anomalous volume dependence in $p\cot\delta$. This indicates that breakup (inelastic) effects are numerically negligible for the particular states and volumes analyzed here, and supports the use of the finite-volume dineutron approximation as an intermediate step toward $2n2n$ scattering.
Table 1: Ratio of neutron-dineutron to neutron-neutron scattering lengths for different dineutron binding energies $B_{2n}$ at a lattice spacing $a=1.64$ fm.
$B_{2n}$ $a_{n-2n}/a_{n-n}$ error
$0.01$ $3.35\times 10^{-7}$ $2.70$
$0.03$ $1.29$ $0.99$
$0.05$ $1.13$ $0.27$
$0.07$ $1.09$ $0.11$
$0.09$ $1.08$ $0.06$
We now turn to the extraction of the $S$-wave 2n2n-$2n$ phase shifts from our finite-volume spectra. For the realistic N3LO chiral EFT Hamiltonian of Ref. Elhatisari and others (2024), the infinite-volume two-neutron system is not bound, but it corresponds to a virtual state. Nevertheless, in the periodic boxes used here the lowest nnnn $S$-wave level is shifted below the two-neutron threshold and can be treated as a shallow finite-volume dimer (a quasi-bound dineutron). This provides a practical intermediate description in which the four-neutron finite-volume levels are interpreted as two composite dimers in relative motion, and Lüscher’s relation can be used to map the discrete energies to elastic 2n2n-$2n$ phase shifts.
For each box size $L$ we compute the four-neutron energy $E_{4n}^{(L)}$ and the two-neutron energy $E_{2n}^{(L)}$ in the same periodic volume. The latter defines the effective finite-volume “binding” of the quasi-bound $2n$ level, $B_{2n}^{(L)}\equiv-E_{2n}^{(L)}$. To extract the relative momentum $p$ of the two-dineutron system we invert the finite-volume dispersion relation for composite dineutrons, Eq. (6). This inversion requires as input the infinite-volume dineutron binding energy $B_{2n}$. For the N3LO interaction this scale tends to zero as $L\!\to\!\infty$ and cannot be determined directly from our largest volumes. We therefore treat $B_{2n}$ as an auxiliary near-threshold parameter and scan representative values $B_{2n}=0.01$$0.09$ MeV (all below the smallest $B_{2n}^{(L)}$ encountered at our largest $L$), using the resulting spread as an estimate of the residual systematic uncertainty associated with employing the finite-volume dineutron picture when the underlying $nn$ pole is virtual.
Figure 3: Dineutron-dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation in Eq. (6), for the N3LO chiral EFT interaction of Ref. Elhatisari and others (2024). Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
The resulting 2n2n-$2n$ $S$-wave phase shifts are shown in Fig. 3. At the smallest extracted relative momenta (largest boxes), the phase shifts show the largest sensitivity to the auxiliary near-threshold input $B_{2n}$ used in the inversion of Eq. (6). While some of the $B_{2n}$ values yield slightly negative $\delta(p)$ in this regime, the low-momentum trend systematically bends toward $\delta(p)\to 0$ as $B_{2n}\to 0$, as expected for short-range interactions. We therefore do not draw a firm conclusion from the sign of the lowest-momentum points in Fig. 3. As the momentum increases, $\delta(p)$ rises, becomes positive, and exhibits a shallow maximum of order $10^{\circ}$ around $p\simeq 60$ MeV, before decreasing again and turning negative at the highest momenta shown. The dependence on the scanned $B_{2n}$ values is mild over most of the range and is most pronounced only at the smallest momenta, where the dineutron size is largest and finite-volume effects are maximally amplified.
In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ displays a clear non-monotonic structure, indicating a weak attractive feature. However, we do not observe the characteristic rapid rise through $90^{\circ}$ that would signal a narrow elastic resonance. The shaded band in Fig. 3 highlights $p\simeq 60$$84$ MeV, which corresponds to $L\simeq 20$$15$ fm and (for the same volumes) to confined $4n$ energies in the range $E_{4n}\simeq 1.7$$3.3$ MeV (cf. Fig. 1), overlapping the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.
IV Summary
We have investigated the tetraneutron system confined in a cubic box using nuclear lattice effective field theory, with box sizes up to $L=30$ fm. Using the N3LO chiral EFT interaction, the confined $4n$ energy decreases gradually with increasing $L$ and remains concave over the explored volumes, showing no plateau that would be characteristic of a narrow resonance. For comparison, we also repeat the confined-spectrum calculation with an SU(4)-symmetric interaction fitted to the 1S0 phase shifts; the SU(4) results are consistent with the N3LO trend and are reported in 1.
To interpret possible $4n$ correlations in terms of two-dineutron dynamics, we employ the finite-volume dineutron approximation, where the lowest nnnn $S$-wave level in a periodic box is quasi-bound even though the infinite-volume $nn$ system corresponds to a virtual state. Within this framework, we validate the Lüscher-based extraction by studying $n$-$n$ and nn-$2n$ scattering. The ratio of scattering lengths $a_{n\text{-}2n}/a_{n\text{-}n}$ exhibits the expected universal behavior for near-threshold systems, with deviations only when $B_{2n}\approx 0.01$ MeV, where the proximity to the virtual-state regime amplifies finite-volume and lattice-spacing effects.
We then extract the 2n2n-$2n$ $S$-wave phase shifts from finite-volume $4n$ and $2n$ energies using Lüscher’s method together with the composite-dineutron finite-volume relation. At small momenta, the negative phase shifts indicate effective repulsion between dineutrons, consistent with effective field theory expectations near unitarity. In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ becomes positive and shows a shallow non-monotonic structure, with a maximum of order $10^{\circ}$ around $p\simeq 60$$84$ MeV, but without a rapid rise through $90^{\circ}$ typical of a narrow elastic resonance. This momentum band corresponds to box sizes $L\simeq 20$$15$ fm and confined $4n$ energies of 1.71.73.33.3 MeV, which lie close to the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.
V Acknowledgments
We are grateful for discussions with Kouichi Hagino, Hiroyuki Sagawa, Furong Xu, and the members of the NLEFT Collaboration. This work is supported by “the Fundamental Research Funds for the Central Universities”, National Natural Science Foundation of China under Grant No. U2541242 and 12435007. The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of UGM was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of YK was supported in part by the Institute for Basic Science (IBS-R031-D1). LW gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University, National Supercomputing Center of Korea with supercomputing resources including technical support (KSC-2024-CHA-0001, KSC-2025-CHA-0004). SS and UGM gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).
Supplementary Material
V.1 Energy extrapolation
For each $L$, the energy is computed at several projection times $\tau$ and extrapolated to the infinite-time limit $E_{\infty}^{(L)}$ using Shen et al. (2025)
$E^{(L)}(\tau)=\frac{E_{\infty}^{(L)}+\left(E_{\infty}^{(L)}+d^{(L)}\right)c^{(L)}e^{-d^{(L)}\tau}}{1+c^{(L)}e^{-d^{(L)}\tau}}$ (S1)
where $E_{\infty}^{(L)}$, $d^{(L)}$, and $c^{(L)}$ are fitting parameters. Details of the extrapolation procedure are provided in 1. We find that the dominant contribution to the interaction energy comes from the 1S0 channel, consistent with previous studies Lazauskas and Carbonell (2005). Other contributions, including the three-nucleon force, are negligibly smaller. For comparison, we also employ an SU(4) symmetric interaction Lu et al. (2019) fitted to the 1S0 phase shift, with the form:
$V_{\rm SU(4)}=\frac{1}{2}C_{2}\sum_{\mathbf{n}}\tilde{\rho}(\mathbf{n})^{2}.$ (S2)
The density operator is defined as
$\tilde{\rho}(\mathbf{n})=\sum_{i}\tilde{a}_{i}^{\dagger}(\mathbf{n})\tilde{a}_{i}(\mathbf{n})+s_{\rm L}\sum_{|\mathbf{n}^{\prime}-\mathbf{n}|=1}\sum_{i}\tilde{a}_{i}^{\dagger}(\mathbf{n}^{\prime})\tilde{a}_{i}(\mathbf{n}^{\prime}),$ (S3)
where $i$ denotes the spin index, $s_{\rm L}$ is a local smearing parameter, and $\tilde{a}_{i}^{\dagger}(\mathbf{n})$ represents the nonlocally smeared creation operator Elhatisari and others (2016):
$\tilde{a}_{i}^{\dagger}(\mathbf{n})={a}_{i}^{\dagger}(\mathbf{n})+s_{\rm NL}\sum_{|\mathbf{n}^{\prime}-\mathbf{n}|=1}{a}_{i}^{\dagger}(\mathbf{n}^{\prime}).$ (S4)
Given the dilute nature of the system, we calculate this SU(4) interaction at a larger lattice spacing $a=1.64$ fm for comparison. The fitted parameters are $s_{\rm NL}=0.1$, $s_{\rm L}=7.3918\times 10^{-2}$, and $C_{2}=-8.4272\times 10^{-6}\,$MeV-2.
By employing Monte Carlo sampling to calculate the energies of the tetraneutron system at different values of $\tau$ for each lattice size $L$, and using Eq. (S1) to extrapolate to infinite projection time, the extrapolation results for both the N3LO and SU(4) interactions are obtained, as shown in Fig. S1 and Fig. S2, respectively.
Figure S1: The energy extrapolation of the four-neutron system with the N3LO interaction at various lattice sizes $L$.
Figure S2: The energy extrapolation of the four-neutron system with the SU(4) interaction at various lattice sizes $L$.
V.2 Energy derivative
Following Refs. Maier et al. (1980); Zhang et al. (2008), we compute the derivatives of the energy with respect to the box size $L$, as shown in Fig. S3. The numerical derivatives, particularly the second derivative for the N3LO interaction, exhibit relatively large uncertainties due to Monte Carlo sampling and projection-time extrapolation. For comparison, we also show results smoothed using the Savitzky-Golay filter Savitzky and Golay (1964). In neither case do we observe the clear signature of a resonance, which would require $\partial^{2}E/\partial L^{2}$ to cross zero. The energy curve remains concave and does not transition to convex behavior.
Figure S3: Energy, energy gradient, and second derivative of the four-neutron system for the N3LO and SU(4) interactions as functions of the lattice size $L$.
V.3 finite-volume extrapolation for dineutron
Figure S4 shows the finite-volume dependence of the two-neutron ground-state energy computed with the SU(4) interaction. To assess the infinite-volume limit we use an empirical exponential form, $E(L)=E_{\infty}+c\,e^{-\delta E\,L}/L$, and systematic large-$L$ expansions in $1/L$. The latter follow from the threshold expansion of Lüscher’s finite-volume quantization condition for short-range interactions, which predicts power-law finite-volume shifts of the lowest two-body level (with leading behavior $\propto 1/L^{3}$Luscher (1986, 1991); Beane et al. (2006). Including higher order terms yields an infinite-volume energy consistent with zero, $E_{\infty}\approx 0$.
Figure S4: Dineutron energy as a function of the cubic box length $L$ for the SU(4) interaction. The curves correspond to infinite-volume extrapolations using different numbers of terms in the fit function.
V.4 Phase shift
The neutron-neutron 1S0 scattering phase shift $\delta$, extracted from finite-volume two-neutron energies via Ls̈cher’s method, is shown in Fig. S5. Results obtained with the SU(4) and N3LO interactions are compared with experimental data Stoks et al. (1993).
Figure S5: Neutron-neutron 1S0 scattering phase shift $\delta$ as a function of relative momentum pp. The points are calculated from finite-volume energies using Lüscher’s method for the SU(4) and N3LO interactions, and are compared to the experimental analysis of Ref. Stoks et al. (1993).
To examine whether the dineutron approximation can be used to study dineutron systems and scattering properties, we performed calculations of phase shifts for the simpler neutron-neutron and neutron-dineutron systems. The elastic two-body scattering phase shift is expanded via the effective range expansion as
$p\cot\delta=-\frac{1}{a_{n-n}}+\frac{1}{2}r_{n-n}p^{2}+\mathcal{O}(p^{4}),$ (S5)
where $p$ is the relative momentum of the two bodies, $a_{n-n}$ is the neutron-neutron scattering length, and $r_{n-n}$ is the neutron-neutron effective range. Using Eq. (S5) for extrapolation, the neutron-dineutron scattering length $a_{n-2n}$ and effective range $r_{n-2n}$ can also be determined by fitting the momentum and phase shift data at different $L$.
To test the universality of the dineutron approximation under pure contact interactions, we varied the dineutron binding energy from 0.1 MeV down to nearly zero at lattice spacings $a=1.32,\;1.64,\;1.97\;\text{fm}$, computed the relation between the dimensionless quantity $ap\cot\delta$ and $ap$ as shown in Fig. S6. Except for the system with a binding energy of 0.01 MeV, all results align with the expectations, indicating the validity of the dineutron approximation.
Based on Eq. (S5), we calculated the neutron-dineutron scattering length $a_{n-2n}$, neutron-neutron scattering length $a_{n-n}$, and their ratio $a_{n-2n}/a_{n-n}$ for different lattice spacings corresponding to a dineutron binding energy $B_{2n}=0.01~\text{ MeV}$. The results are listed in Table S1. Here, the values of $a_{n-n}$ are large and negative across all lattice spacings, indicating a virtual state in the system, while the neutron-dineutron system also exhibits large negative scattering lengths. Additionally, Table S2 presents the ratio of neutron-dineutron to neutron-neutron scattering lengths for dineutron binding energies ranging from $B_{2n}=0.01$ to $0.1~\text{ MeV}$ at different lattice spacings. The results show that for lattice spacings $a=1.32,1.64,$ and $1.97~\text{ fm}$, the values of $a_{n-2n}/a_{n-n}$ exhibiting a universal behavior independent of the lattice spacing.
Figure S6: Neutron-neutron (a-c) and neutron-dineutron (d-f) scattering phase shifts calculated using Lüscher’s finite-volume method with the SU(4) interaction at lattice spacings $a=1.32$ fm (a,d), $1.64$ fm (b,e), and $1.97$ fm (c,f). Points represent lattice data for different $B_{2n}$ values, and lines show fits based on the effective range expansion.
Table S1: Scattering lengths and errors for different values of $a$ values ($B_{2n}=0.01$ MeV).
quantity $a=1.32$ fm $a=1.64$ fm $a=1.97$ fm
value error value error value error
$a_{n-n}$ -418.952822 4.714405 $3.625460\times 10^{8}$ $1.057483\times 10^{14}$ -433.063197 6.848189
$a_{n-2n}$ -37.524722 0.948073 -67.299539 2.624853 -44.818149 0.847960
$a_{n-2n}/a_{n-n}$ 0.089568 0.002477 $1.856303\times 10^{-7}$ 0.054145 0.103491 0.002552
Table S2: The ratio of neutron-dineutron to neutron-neutron scattering lengths for different lattice spacings and dineutron binding energies.
$B_{2n}$ $a=1.32$ fm $a=1.64$ fm $a=1.97$ fm
$a_{n-2n}/a_{n-n}$ error $a_{n-2n}/a_{n-n}$ error $a_{n-2n}/a_{n-n}$ error
$0.01$ $0.089568$ $0.002477$ $1.86\times 10^{-7}$ $0.054145$ $0.103491$ $0.002552$
$0.02$ $4.370859$ $0.739432$ $1.955848$ $0.104053$ $2.072985$ $0.093956$
$0.03$ $1.547733$ $0.054318$ $1.318567$ $0.028219$ $1.276168$ $0.018059$
$0.04$ $1.293470$ $0.027550$ $1.190182$ $0.015986$ $1.155832$ $0.009129$
$0.05$ $1.200866$ $0.018532$ $1.138108$ $0.010687$ $1.113122$ $0.005522$
$0.06$ $1.154382$ $0.013838$ $1.111648$ $0.007609$ $1.094222$ $0.003659$
$0.07$ $1.127364$ $0.010883$ $1.096796$ $0.005587$ $1.085393$ $0.002799$
$0.08$ $1.110502$ $0.008832$ $1.088154$ $0.004205$ $1.081597$ $0.002647$
$0.09$ $1.099464$ $0.007320$ $1.083139$ $0.003294$ $1.080561$ $0.002883$
$0.10$ $1.092066$ $0.006170$ $1.080411$ $0.002781$ $1.081122$ $0.003255$
Figure S7: Dineutron-dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation in Eq. (6), for the SU(4) interaction. Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
We also calculated dineutron-dineutron $S$-wave scattering phase shifts for the SU(4) interaction, as shown in Fig. S7. In the momentum range 60-120 MeV, the phase shift also shows some attraction. Similar to the case with the N3LO interaction, no rapid rise of the phase shift to 90° is observed. The shaded region corresponds to $p\simeq$ 60–80 MeV, which for a box size of $L\simeq$ 20-15 fm yields a calculated tetraneutron energy range of $E\simeq 1.7$–3.0 MeV.